OpenStudy (pulsified333):

Find the derivative of y=−13e^(−0.9t)

OpenStudy (pulsified333):

@satellite73

zepdrix (zepdrix):

Remember your exponential derivative?\[\large\rm \frac{d}{dx}e^{stuff}=e^{stuff}\frac{d}{dx}stuff\]We get the same thing back, but with chain rule.

zepdrix (zepdrix):

The -13 is a constant coefficient, it has no effect on the differentiation process,\[\large\rm \frac{d}{dx}\left(-13e^{-0.9t}\right)=-13\frac{d}{dx}e^{-0.9t}\]

zepdrix (zepdrix):

\[\large\rm =-13e^{-0.9t}\cdot(-0.9t)'\]So just chain rule :) Multiply by derivative of the exponent.

OpenStudy (pulsified333):

oh okay

OpenStudy (pulsified333):

@zepdrix that didn't work

zepdrix (zepdrix):

Do you see the ( )' at the end? The tick mark is to indicate that you need to take the derivative of that portion.

zepdrix (zepdrix):

Do you understand how to take derivative of 0.9t?

OpenStudy (pulsified333):

9/10?

zepdrix (zepdrix):

Good good good. I should have said -0.9t, which turns into -0.9 through differentiation.

zepdrix (zepdrix):

Simplify the problem by multiplying the -13 and -0.9 together.

OpenStudy (pulsified333):

ah now it worked :D thanks

zepdrix (zepdrix):

yay :)